# CRAFT: Training-Free Cascaded Retrieval for Tabular QA

### v1

@@ Abstract
Table Question Answering (TQA) involves retrieving relevant tables from a large corpus to answer natural language queries. Traditional dense retrieval models, such as DTR and ColBERT, not only incur high computational costs for large-scale retrieval tasks but also require retraining or fine-tuning on new datasets, limiting their adaptability to evolving domains and knowledge. In this work, we propose CRAFT, a cascaded retrieval approach that first uses a sparse retrieval model to filter a subset of candidate tables before applying more computationally expensive dense models and neural re-rankers. Our approach achieves better retrieval performance than state-of-the-art (SOTA) sparse, dense, and hybrid retrievers. We further enhance table representations by generating table descriptions and titles using Gemini Flash 1.5 Team et al. (2025). End-to-end TQA results using various Large Language Models (LLMs) on NQ-Tables Herzig et al. (2021), a subset of the Natural Questions Dataset, demonstrate CRAFT effectiveness.
@@
$**{}^{\textbf{*}}$darsh $**{}^{\textbf{*}}$ushal Raj Bhandari2  Jianxi Gao2 Soham Dan3Vivek Gupta1 1Arizona State University 2Rensselaer Polytechnic Institute 3Microsoft asing725@asu.edubhandk@rpi.edu
@@ 1 Introduction
Recent work has started to tackle open-domain TQA by developing specialized retrieval models. Early approaches adopted sparse or dense text retrievers on flattened tables (e.g., BM25 or bi-encoder models)Fan et al. (2024); Robertson and Zaragoza (2009). For example, the Dense Table Retriever (DTR) introduced by Herzig et al. (2021) augments a bi-encoder with row and column features to better capture table structure. More recent methods explicitly incorporate fine-grained structure: Jin et al. (2023) propose a syntax- and structure-aware retriever that represents each question phrase and each table header and values separately, achieving state-of-the-art results on the NQ-Tables benchmark. Likewise, Li et al. (2025) introduces THYME, a field-aware hybrid retriever that tailors matching strategies to different table fields (e.g., cell values vs. captions). THYME significantly outperforms prior baselines on open-domain table retrieval, highlighting the gains from table-specific modeling. Similarly, Pan et al. (2022) introduced T-RAG, an end-to-end table QA framework that combines table retrievers with sequence-to-sequence generators for retrieval-augmented generation. These advances, however, often rely on complex architectures or extensive fine-tuning on the TQA dataset. This raises an intriguing question: Can off-the-shelf pretrained models, when integrated into a carefully designed retrieval pipeline, achieve competitive performance to expensive domain-specialized, fine-tuned methods? To this end, we propose CRAFT, a modular and scalable multi-stage retrieval framework for TQA that harnesses pretrained models without relying on dataset-specific fine-tuning.
Our work introduces CRAFT, a training free cascaded tabular retriever algorithm that achieves competitive performance with state-of-the-art tabular retrievers requiring extensive in-domain training. Through comprehensive ablation studies, we demonstrate the robustness of CRAFT across varying retrieval threshold $K$, highlighting its adaptability and stability. Building on this foundation, we develop a complete end-to-end TQA system that outperforms all existing TQA-in-the-wild , showcasing the effectiveness of our retrieval approach in real-world, open-domain settings.
2 Methodology
@@
To create a retrieval setup without dataset-specific fine-tuning, we adopt a modular framework that leverages off-the-shelf models with multi-stage retrieval, with more complex models for each successive stage. As shown in Figure 1, CRAFT involves prepossessing followed by stage-wise table retrieval using increasingly expressive models, culminating in answer generation via a pretrained LLM. Our approach demonstrates that pretrained models can effectively retrieve relevant tabular data and generate accurate answers compared to the state-of-the-art framework Zhao et al. (2024). The pipeline consists of leveraging preprocessing followed by cascaded table retrieval and end-to-end answer generation to maximize semantic alignment between the input queries and tabular content using off-the-shelf embeddings models.
@@ 2.1 Preprocessing
The preprocessing step involved three key operations to enhance the query and table representations for effective retrieval. Firstly, Gemini-1.5-Flash Team et al. (2025) was employed to generate sub-questions derived from each original query, thus decomposing complex queries into more granular components. Secondly, table titles and comprehensive descriptions were generated using Gemini-1.5-Flash, enriching contextual understanding and enabling more precise downstream retrieval. Finally, all rows across every table were ranked using a Sentence Transformer (all-mpnet-base-v2) model SentenceTransformers (2022), applied directly to the original queries to assess semantic relevance. The NQ-Tables Dataset statistics are shown in Table 7.
@@ 2.2 Table Retrieval
The table retrieval step involves a rigorous, three-stage filtering methodology to identify highly relevant tables for each query. In the first stage, SPLADE Formal et al. (2021) was applied to a comprehensive data set consisting of 169,898 tables, leveraging generated table titles, column headers, individual cell values, and detailed table descriptions to filter and narrow down the candidate tables to the top 5,000 entries. SPLADE leverages sparse lexical expansion to efficiently score candidate passages using minimal memory, making it well-suited for handling large-scale retrieval tasks.
In the second stage, the Sentence Transformer (all-mpnet-base-v2) provides a fine-grained semantic refinement by representing each table with mini-tables created through appending column headers to each of the top 5 rows. This refinement approach allows the selection of the top 1,000 mini-tables. In the final stage, the text-embedding-3-small OpenAI (2024) model performs a detailed re-ranking on the top 100 mini-tables, carefully assessing these mini-tables based on their semantic embeddings. Concurrently, we enrich all 919 unique queries using sub-question decomposition generated by Gemini-1.5-Flash to maximize their semantic alignment with the selected tables.
@@ 2.3 End-to-End Answer Generation
For the end-to-end table question-answering, we use the final retrieval result on the set of 919 queries with three different pretrained models: Llama3-8B, Qwen2.5-7B, and Mistralv0.2-7B. Each LLM generates answers from prompts that include a varying number of tables ($n=\{1,3,5,8,10\}$). Each prompt contains the table title, column headers, and mini-tables (top 5 rows) for $n$ tables, offering comprehensive yet concise context to support accurate answer generation. In addition, we also experiment with the inclusion of a few-shot prompt to guide the text generation process.
We evaluate model performance and accuracy by computing the F1 score, which quantifies the overlap between generated answers and reference answers, enabling comparative analysis across different models and table settings.
@@
        + Stage 3 41.13 87.16 96.84
CRAFT 41.13 87.16 96.84
Table 1: Table Retrieval Performance on NQ-Tables. R denotes Recall Score. Best performances are in Bold. All models except CRAFT are trained on NQ-Tables. ‘*’ denotes results reported for the full test-set with all 966 queries.
3 Results and Analysis
CRAFT Retrieval Performance.
Table 1 presents the retrieval performance of three groups of methods: Sparse, Dense, and Hybrid Retrievers on the NQ-Tables dataset. Hybrid retrievers show the best overall performance, and dense retrievers outperform Sparse Retrievers. Our method achieves performance comparable to SOTA fine-tuned retrievers, even outperforming them at R@10 and R@50. THYME achieves the best average performance among all retrievers. Compared to SPLADE and DTR our three-stage cascade improves retrieval step by step. In Stage 1, SPLADE finds almost all relevant tables at higher ranks (R@50 = 91.62%) but only 34.38% at the top rank (R@1). In Stage 2, the Sentence-Transformer reranker boosts mid-level performance, increasing R@10 by over 10 percentage points and pushing R@50 above 96%, while still keeping strong overall recall. Finally, the neural reranker in Stage 3 focuses on improving top results, raising R@1 by another 4.5 points and refining R@10 to 87.16%. Although CRAFT achieved a lower recall score at R@1, we observed that for some queries, there are tables in the corpus, other than the gold table, that also contains the answer to the query.
Robustness of CRAFT Retrieval.
@@ 3 Results and Analysis
        + Stage 3 41.02 86.83 96.08 - 0.04
Table 2: Performance of DTR vs CRAFT on perturbed queries. M and S denote Medium and Small model sizes respectively. $\Delta$ is the Avg. change in Recall relative to the original query scores, positive values indicate improvement, and negative values indicate degradation. HN represents a model trained with Hard Negatives.
@@
THYME 35.48 36.14 37.28 37.59 39.16 40.28 39.29 39.41 41.20
CRAFT 39.65 39.29 33.33 46.04 44.06 37.70 45.09 45.98 41.47
CRAFT (5-fewshot) 41.17 41.72 40.36 48.50 39.39 43.65 47.70 39.27 45.81
Table 3: Performance comparison (F1) across different retrievers, LLMs, and values of $n$ with NQ-Tables. The evaluated models include Mistral-7B, Llama3-8B, and Qwen2.5-7B, corresponding to Mistral, Llama3, and Qwen, respectively. Bolded values indicate top results.
To assess robustness under query paraphrasing, we generated perturbed versions of each question using Gemini 2.5 Flash Team et al. (2025), while keeping the ground-truth answer unchanged. Table 2 reports recall scores for DTR, trained on NQ-Tables with and without hard negatives, and our three-stage CRAFT pipeline. We directly use publicly released fine-tuned DTR models for this experiment. With the original queries, CRAFT already outperforms the DTR baselines. When using the perturbed queries, both DTR Medium and Small suffer substantial average drops of –8.38 and –11.82 points respectively. The DPR version trained with hard negatives reduces but does not eliminate this degradation (avg = –5.80 for DTR (M) and –6.64 for DTR (S)). By contrast, CRAFT’s full three-stage pipeline nearly fully recovers its original performance with a slight decrease under paraphrasing.
Ablations Analysis.
Ablation over stages further shows incremental gains: Stage 1 drops performance under perturbation by 3.4 %, Stage 2 partly recovers it, and Stage 3 restores the original recall, as shown in Appendix A.2. These results highlight CRAFT’s superior robustness to query rephrasing compared to DTR.
End-to-End Results.
We evaluate the end-to-end performance of CRAFT across multiple pretrained language models and retrieval configurations. As shown in Table 3, CRAFT consistently outperforms several state-of-the-art baselines across different models: Mistral Jiang et al. (2023), LLaMA3 Grattafiori et al. (2024), and Qwen 2.5 Qwen et al. (2025), for varying numbers of retrieved tables ($n={1,3,5}$). These results highlight the robustness of our modular, training-free approach.
When considering only one retrieved table ($n=1$), CRAFT notably outperforms baseline methods, significantly improving accuracy scores using all the pretrained models. Moreover, incorporating a 5-shot prompt further enhances performance substantially across all models and top-k tables. These results highlight that structured, cascaded retrieval by itself outperforms systems that rely on dataset‑specific training, and that augmenting this retrieval with lightweight few‑shot prompting delivers consistent additional gains. We further explore the impact of performance using different few-shot prompting and the effect on context reduction with CRAFT in Section B.
@@
DPR-RAGE Lin et al. (2023) - 49.68
LI-RAGE Lin et al. (2023) - 54.17
TABERT
   - Mistral-Small-Instruct 50.26 57.14
   - Llama-3.1-70B-Instruct 49.75 56.94
@@
Table 4: F1 scores of CRAFT with larger LLMs compared to dataset-specific baselines for $n=1$ and $n=5$ retrieved tables. Bold indicates the best performance.
@@
Table 4 further emphasizes the effectiveness of CRAFT, particularly when employing larger, instruction-tuned LLMs. Compared to prior dataset-specific baselines, including RAG Lewis et al. (2020), DPR-RAGE, and LI-RAGE Lin et al. (2023). CRAFT consistently delivers superior F1 scores without requiring any dataset-specific training or finetuning. We find that using just a single retrieved table ($n=1$), Mistral-Small-Instruct achieves an F1 score of $50.26$, already surpassing the best-performing baseline, LI-RAGE ($54.17$), even though LI-RAGE benefits from deeper retrieval ($n=5$). With increased retrieval depth ($n=5$), Mistral-Small-Instruct further improves to $57.14$, highlighting significant gains from additional context. Similarly, Llama-3.1-70B-Instruct achieves a notable score of $56.94$ at $n=5$, clearly outperforming current SOTA. These results underscore that CRAFT’s structured, training-free retrieval methodology not only outperforms established baselines but also scales effectively with increased retrieval depth and few-shot prompting. This highlights the strength of CRAFT’s modular design in enabling adaptable, efficient, and high-performing TQA without needing any dataset-specific fine-tuning.
4 Conclusion
We present CRAFT, a modular, training-free, multi-stage cascaded retrieval framework for TQA that matches or exceeds the performance of state-of-the-art systems trained on NQ-Tables. Without any dataset-specific training, it achieves competitive SOTA recall scores, maintains robustness under query perturbation, and delivers a high end-to-end performance score when paired with existing pretrained LLMs. By decoupling retrieval from domain-specific fine-tuning, CRAFT offers a scalable, adaptable solution for dynamic tabular domains where rapid deployment and data scarcity are critical issues.
5 Limitations
While our evaluation focuses on the NQ-Tables benchmark, future validation on additional TQA datasets would help further assess the generalizability of our approach. Although our multi-stage retrieval pipeline, combining sparse filtering, dense semantic reranking, and neural reranking, introduces added inference steps, this design enables significantly improved retrieval accuracy, offering a trade-off between performance and runtime that can be optimized for specific deployment needs.
Appendix A CRAFT: Retrieval Analysis
A.1 Stage-wise Results

### v2

@@ Abstract
Open-Domain Table Question Answering (TQA) involves retrieving relevant tables from a large corpus to answer natural language queries. Traditional dense retrieval models, such as DTR and DPR, not only incur high computational costs for large-scale retrieval tasks but also require retraining or fine-tuning on new datasets, limiting their adaptability to evolving domains and knowledge. In this work, we propose CRAFT, a zero-shot, cascaded retrieval approach that first uses a sparse retrieval model to filter a subset of candidate tables before applying more computationally expensive dense models as re-rankers. To improve retrieval quality, we enrich table representations with descriptive titles and summaries generated by Gemini Flash 1.5, enabling richer semantic matching between queries and tabular structures. Our method outperforms state-of-the-art (SOTA) sparse, dense, and hybrid retrievers on the NQ-Tables dataset. It also demonstrates strong zero-shot performance on the more challenging OTT-QA benchmark, achieving competitive results at higher recall thresholds, where the task requires multi-hop reasoning across both textual passages and relational tables. This work establishes a scalable and adaptable paradigm for table retrieval, bridging the gap between fine-tuned architectures and lightweight, plug-and-play retrieval systems. Code and data are available at: https://coral-lab-asu.github.io/CRAFT/
@@
[…]
@@ 1 Introduction
Recent work has started to tackle open-domain TQA by developing specialized retrieval models. Early approaches adopted sparse or dense text retrievers on flattened tables (e.g., BM25 or bi-encoder models) Fan et al. (2024); Robertson and Zaragoza (2009). The Dense Table Retriever (DTR) introduced by Herzig et al. (2021) augments a bi-encoder with row and column features to better capture table structure. Several methods employ multi-stage cascades to prune candidates before applying costly scoring, with each stage acting as a separate ranking component. For example,  Wang et al. (2011) introduced a two-stage document ranking cascade composed of a candidate generation step and a learned re-ranker, which are trained jointly as a single model. In a similar vein, the ”Retrieve, Read, Rerank” framework from  Hu et al. (2019) embeds retrieval, answer extraction, and re-ranking within a single shared Transformer network that is fine-tuned end-to-end. The Re²G system  Glass et al. (2022) implements a retrieve-rerank-generate pipeline where the re-ranking and generation stages are merged into one BART-based sequence-to-sequence model fine-tuned with knowledge distillation.
@@
More recent methods explicitly incorporate fine-grained structure. Jin et al. (2023) proposed a syntax and structure-aware retriever that represents each question phrase and each table header and values separately, achieving state-of-the-art results on the NQ-Tables benchmark. Likewise, Li et al. (2025) introduces THYME, a field-aware hybrid retriever that tailors matching strategies to different table fields (e.g., cell values vs. captions). It significantly outperforms prior baselines on open-domain table retrieval, highlighting the gains from table-specific modeling. Similarly, Pan et al. (2022) introduced T-RAG, an end-to-end table QA framework that combines table retrievers with sequence-to-sequence generators for retrieval-augmented generation.
These advances, however, often rely on complex architectures or extensive fine-tuning on the TQA dataset. This raises an intriguing question: Can off-the-shelf pretrained models, when integrated into a carefully designed retrieval pipeline, achieve competitive performance to expensive domain-specialized, fine-tuned methods? To this end, we propose CRAFT, a modular and scalable multi-stage retrieval framework for TQA that harnesses pretrained models without relying on dataset-specific fine-tuning.
Our contributions are as follows :
  • We design a modular, cascaded architecture that strategically combines complementary retrieval signals (lexical and semantic) to maximize recall while maintaining efficiency.
  • Through comprehensive ablation studies, we demonstrate the robustness of CRAFT across varying retrieval thresholds $N$, highlighting its adaptability and stability.
  • CRAFT achieves state-of-the-art retrieval recall and end-to-end performance on NQ-Tables, and strong zero-shot performance on OTT-QA, with particularly competitive results at higher recall thresholds, demonstrating the effectiveness of training-free retrieval in real-world, open-domain settings.
  • A mini-table centric cascaded design that reduces token footprint and embedding overhead, achieving up to $33\times$ fewer online embedding calls and $70\%$ shorter contexts. It improves retrieval and downstream QA efficiency without compromising accuracy.
2 Retrieval Benchmarks
We used two datasets (a.) NQ-Tables Herzig et al. (2021), and (b.) OTT-QA Chen et al. (2020) to evaluate our CRAFT approach. Both datasets are paired with extensive table corpora. NQ-Tables focuses on single-hop factual queries, whereas OTT-QA requires more intricate multi-hop reasoning across texts and tables. We use the subset relevant to the table for retrieval evaluation. Each query in both datasets is associated with approximately one gold table, ensuring consistency in supervision and facilitating direct comparison of retrieval performance. Table 1 provides an overview of the NQ-Tables and OTT-QA benchmarks.
NQ-Tables OTT-QA
Train Test Train Test
Total Query Count 9,594 966 41,469 2,214
Unique Query Count 9,594 919 41,469 2,214
Avg. # Words 8.94 8.90 21.79 22.82
Table Count 169,898 419,183
Avg. # Rows 10.70 12.90
Avg. # Columns 6.10 4.80
# Golden Tables / Query 1.00 1.05 1.00 1.00
Table 1: Statistics of the NQ-TABLES and OTT-QA benchmarks.
Together, these statistics underline the scalability and diversity of the evaluation setting, where NQ-Tables offer a more controlled environment and OTT-QA presents a challenging, open-domain scenario that better tests model generalization.
3 Our CRAFT Approach
@@
To create a retrieval setup without dataset-specific fine-tuning, we adopt a modular framework that leverages off-the-shelf models with multi-stage retrieval, with more complex models for each successive stage. As shown in Figure 1, CRAFT involves pre-processing followed by stage-wise table retrieval using increasingly expressive models, culminating in answer generation via a pretrained LLM. Our approach demonstrates that pretrained models can effectively retrieve relevant tabular data and generate accurate answers. The pipeline consists of leveraging pre-processing followed by cascaded table retrieval and end-to-end answer generation to maximize semantic alignment between input queries and tabular content using off-the-shelf embedding models.
@@ 3.1 Preprocessing
The preprocessing pipeline comprised three key operations designed to enhance both query and table representations for effective retrieval. Firstly, Gemini-1.5-Flash Team et al. (2025) was employed to generate sub-questions from each original query and to produce informative table titles and detailed descriptions, thereby improving query decomposition and contextual understanding. Finally, all table rows were ranked according to their semantic relevance to the queries using model (all-mpnet-base-v2) from Sentence Transformer Reimers and Gurevych (2019) for the NQ-Tables dataset, and Jina Embeddings v3 Sturua et al. (2024) for the OTT-QA dataset.
@@ 3.2 Table Retrieval
The table retrieval step follows a unified three-stage pipeline applied to NQ-Tables and OTT-QA, with dataset-specific inputs and model choices. In the first stage, SPLADE Formal et al. (2021) , a pre-trained sparse lexical-expansion retriever, was applied to a comprehensive dataset consisting of 169K and 419K Tables in NQ-Tables and OTT-QA dataset, respectively, leveraging generated table titles, column headers, individual cell values, and table descriptions to filter and narrow down the candidate tables to the top 5,000 entries. SPLADE leverages sparse lexical expansion to efficiently score candidate passages using minimal memory, making it well-suited for handling large-scale retrieval tasks.
In the Second Stage, mini-tables are constructed by appending column headers to each of the top 5 rows and then scoring these mini-tables to produce a compact shortlist. For NQ-Table, the Sentence Transformer all-mpnet-base-v2, a 768 dimensional pre-trained bi-encoder model optimized for semantic similarity with solid zero/few-shot retrieval performance across heterogeneous text, is used for fine-grained semantic matching due to strong sentence-level retrieval quality and robust generalization. For OTT-QA, Jina Embeddings v3, a multilingual long-context embedding model with task adapters that improve robustness on mixed schema/numeric text, is used to better handle longer row-level contexts and numeric content typical of diverse table formats in OTT-QA. This stage selects the top K mini-tables for final reranking.
Lastly, Stage three, the top-$K$ mini-tables are re-ranked using state-of-the-art embedding models optimized for semantic precision and computational efficiency. For the NQ-Table dataset, we employ text-embedding-3-small OpenAI (2024) and text-embedding-3-large, for efficient semantic re-ranking. For the OTT-QA dataset, re-ranking is performed using gemini-embedding-001 Lee et al. (2025), designed for long-context understanding and production-grade scalability. The Gemini embedding space enables robust semantic scoring without additional query or document augmentation, preserving the integrity of dense representations and preventing noise amplification.
Overall, this cascaded retrieval framework progressively transitions from high-recall sparse retrieval to high-precision semantic reranking, achieving efficient large-scale performance while minimizing information loss between stages.
@@ 3.3 End-to-End Answer Generation
For the end-to-end table question-answering, we use the final retrieval result from CRAFT Stage 3, using three different pre-trained models: Llama3-8B Grattafiori et al. (2024), Qwen2.5-7B Qwen et al. (2025), and Mistral-7B Jiang et al. (2023). Each LLM generates answers from prompts that include a varying number of tables ($n=\{1,3,5,8,10\}$). Each prompt contains the table title, column headers, and mini-tables (top 5 rows) for $n$ tables, offering a comprehensive yet concise context to support accurate answer generation. Additionally, we experimented with the inclusion of a few-shot prompt to guide the text generation process. We evaluate model performance and accuracy by computing the F1 score, which quantifies the overlap between the generated answers and the reference answers, enabling a comparative analysis across different models and table settings.
4 Results and Analysis
4.1 Retrieval Performance.
Table 2 and Table 3 present the retrieval performance of three categories of methods Sparse, Dense, and Hybrid retrievers on the NQ-Tables and OTT-QA datasets, respectively. We conduct a comprehensive comparison across traditional sparse models such as BM25 and SPLADE, dense retrievers including TaPas Herzig et al. (2020), DPR Karpukhin et al. (2020), BIBERT  Yates et al. (2021), DTR, T-RAG, and SSDR, as well as hybrid approaches like Re2G, THYME, DHR Lin et al. (2023a), and BIBERT+SPLADE.
@@ 4 Results and Analysis
[…]
CRAFT 49.84 86.83 97.17
Table 2: Table retrieval performance on NQ-Tables. All models except CRAFT are trained on NQ-Tables. : results on the full test set (966 queries). $^{*\text{l}}$ and $^{*\text{s}}$: use text-embedding-3-large and text-embedding-3-small in Stage 3, respectively. § denotes models trained with the THYME setup. denotes the use of the original released model checkpoints. R = Recall. Best results in bold; second-best are underlined.
CRAFT demonstrates strong retrieval performance across both datasets, achieving state-of-the-art results on NQ-Tables and competitive zero-shot generalization on OTT-QA.
Retrieval on NQ Tables.
On NQ-Tables, CRAFT progressively improves through its three-stage cascade: In Stage 1, R@10 is 72.90, and in Stage 2, it is boosted to 82.91, and Stage 3’s neural reranker further refines top-ranked results, raising R@1 to 41.13 and R@10 to 87.16. Using the higher-dimensional text-embedding-3-large model increases R@1 to 49.84.
Retrieval on OTT-QA.
CRAFT maintains competitive recall at higher retrieval depths despite operating entirely without task-specific training. While fine-tuned hybrid retrievers such as THYME and BIBERT+SPLADE achieve superior Recall@1 through dataset-specific optimization, CRAFT approaches their performance at R@10 (89.88 vs. 91.10) and R@50 (96.07 vs. 96.34), with gaps of only 1.22 and 0.27 points respectively. The stage-wise improvements mirror those observed on NQ-Tables: Stage 2 delivers a substantial gain of +21.0 points at R@10 over Stage 1, while Stage 3 contributes an additional +8.0 points. This consistent progression across both datasets demonstrates that CRAFT’s modular, training-free cascade effectively amplifies recall and maintains strong cross-dataset generalization, even when scaling from 169K tables (NQ-Tables) to 419K tables (OTT-QA) with more complex multi-hop reasoning requirements.
Model R@1 R@10 R@50
Sparse Retrieval Models
BM25 23.98 51.94 69.11
SPLADE 62.74 89.52 95.21
Dense Retrieval Models
BIBERT 56.82 86.50 94.26
DPR 53.43 85.95 93.22
TAPAS 57.86 86.77 94.04
DTR 42.10 75.75 88.80
SSDR(im) 56.96 86.22 93.55
Hybrid Retrieval Models
BIBERT+BM25 59.49 86.81 94.67
DHR 63.64 88.48 95.30
BIBERT+SPLADE 64.72 91.01 96.34
THYME 66.67 91.10 96.16
CRAFT Stage-wise
Stage 1 40.74 60.89 88.17
   + Stage 2 47.33 81.88 91.82
    + Stage 3 55.56 89.88 96.07
CRAFT 55.56 89.88 96.07
Table 3: Table Retrieval Performance on OTT-QA. R denotes recall score. Best performances are shown in bold, and second-best are underlined. All models, except CRAFT, are trained on OTT-QA. SPLADE in the table is the supervised baseline, whereas CRAFT Stage 1 uses the base SPLADE model.
4.2 Robustness of CRAFT Retrieval.
To assess robustness under query paraphrasing, we generated perturbed versions of each question using Gemini 2.5 Flash, while keeping the ground-truth answers unchanged. For example:
Original
: Philadelphia is known as the city of what?
Perturbed
: What descriptor follows “the city of” when referencing Philadelphia?
Answer
: Brotherly Love
The prompt used for generating the perturbed queries is detailed in Figure 5.
Analysis.
Table 4 reports recall scores on perturbed queries for DTR, trained on NQ-Tables with and without hard negatives, and our three-stage CRAFT pipeline. We chose it as a baseline because it is a traditional, well-known, and widely adopted model. The checkpoint used in our experiments is publicly available and has been finetuned on the NQ-Tables dataset, making it an ideal baseline for comparing our Cascade Approach against a standard dense embedding method. Additionally, it serves as a common baseline used in many papers that report scores on the NQ-Tables Dataset, like LI-RAGE, DTR, THYME, CLTR, SSDR, and T-RAG, facilitating easier comparison.
@@
   + Stage 3 $^{*\text{s}}$ 41.02 86.83 96.08 - 0.04
Table 4: Performance of DTR vs CRAFT with perturbed queries on NQ-Tables. M and S denote Medium and Small model sizes respectively. $\Delta$ is the average change in Recall relative to the original query scores, positive values indicate improvement, and negative values indicate degradation. HN represents a model trained with Hard Negatives. $^{*\text{s}}$ indicate the use of text-embedding-3-small model in Stage 3.
With the original queries, CRAFT already outperforms the DTR baselines. When using the perturbed queries, both DTR Medium and Small suffer substantial average drops of –8.38 and –11.82 points, respectively. The DTR version trained with hard negatives reduces but does not eliminate this degradation (avg = –5.80 for DTR (M) and –6.64 for DTR (S)). By contrast, CRAFT’s full three-stage pipeline nearly fully recovers its original performance with a slight decrease in paraphrasing.
Stage-wise Analysis.
Analysis over stages further shows incremental gains: Stage 1 drops performance under perturbation by 3.4 %, Stage 2 partly recovers it, and Stage 3 restores the original recall, as shown in Table 4. These results highlight CRAFT’s superior robustness to query rephrasing compared to DTR.
4.3 End-to-End Results.
We evaluate the end-to-end performance of CRAFT across multiple pretrained language models and retrieval configurations. For end-to-end QA, our approach aligns with recent work, such as THYME and EE-BIBERT, which report results on larger open-source models like Llama3-8B, Qwen2.5-7B, and Mistral-7B and closed-source model GPT-4o.
Results on NQ-Tables.
As shown in Table 5, CRAFT consistently outperforms several state-of-the-art baselines across different models, for varying numbers of retrieved tables (n=1,3,5n={1,3,5}). These results highlight the robustness of our modular, training-free approach. Beyond higher recall, mini-table filtering and compact context windows reduce irrelevant content and allow more evidence(more retrieved tables) per query, so even semantically correct non-gold tables can still support higher end-to-end F1.
@@
THYME 35.48 36.14 37.28 37.59 39.16 40.28 37.20 39.29 41.20
CRAFT 39.13 38.31 39.73 45.28 40.76 43.52 44.53 40.55 46.49
Table 5: Performance comparison (F1) across different retrievers, LLMs, and values of $n$ with NQ-Tables. The evaluated models include Mistral-7B-Instruct, Llama3-8B-Instruct, and Qwen2.5-7B-Instruct, corresponding to Mistral, Llama3, and Qwen, respectively. Bolded values indicate top results.
@@
DPR-RAGE Lin et al. (2023b) - 49.68
LI-RAGE Lin et al. (2023b) - 54.17
CRAFT
@@
Table 6: F1 scores of CRAFT on NQ-Tables with larger open source and closed source LLMs compared to dataset-specific baselines for $n=1$ and $n=5$ retrieved tables. Bold indicates the best performance.
Effect of Larger LLMs.
@@
Table 6 further emphasizes the effectiveness of CRAFT, particularly when employing larger, instruction-tuned LLMs. Compared to prior dataset-specific baselines, CRAFT consistently delivers superior F1 scores. We find that using just a single retrieved table ($n=1$), GPT 4o achieves an F1 score of $56.76$, already surpassing the best-performing baseline, LI-RAGE ($54.17$), even though it benefits from deeper retrieval ($n=5$). With increased retrieval depth ($n=5$), GPT-4o further improves to $67.72$, highlighting significant gains from additional context. Similarly, Mistral-Small-Instruct and Llama-3.1-70B-Instruct achieve notable scores of $57.14$ and $56.94$ at $n=5$, clearly outperforming LI-RAGE (current SOTA).
Effect of Few-Shots.
To evaluate the scalability and robustness of CRAFT, under zero-shot and few-shot prompting as the number of retrieved tables increases ($n=1,3,5,8,10$). As shown in Table 14 (Appendix Section C), all models exhibit a clear upward trend in F1 score with deeper retrieval, demonstrating that structured multi-stage retrieval alone substantially narrows the gap with dataset-specific fine-tuning. However, incorporating a small number of in-context examples does not consistently yield significant improvements across models. Moreover, performance gains taper beyond $n=5$, suggesting an optimal retrieval depth that balances information richness with the risk of noise accumulation. This effect becomes more pronounced as both $n$ and the number of few-shot examples increase, leading to greater performance variability.
Result with OTT-QA.
We further provide end-to-end performance on the OTT-QA dataset under an oracle setting, where retrieved tables are paired with the gold passage. Under this oracle setup, where retrieved tables are paired with their corresponding gold passages to isolate the effect of retrieval quality, our end-to-end OTT-QA evaluation shows that Stage 3 (MT) consistently achieves the best balance between recall and precision. As this setup is not directly comparable to prior baselines, we present these results in the Appendix Section C for completeness.
These results underscore that CRAFT’s structured, training-free retrieval methodology not only outperforms established baselines but also scales effectively with increased retrieval depth and few-shot prompting. This highlights the strength of CRAFT’s modular design in enabling adaptable, efficient, and high-performing Table-QA without requiring any dataset-specific fine-tuning.
4.4 Computational Efficiency
Figure 2: End-to-end performance comparison across context lengths with the NQ dataset. The figure illustrates the relationship between average context length and F1 score across three models, evaluated over ten different numbers of retrieved tables ($n$) on the NQ dataset. FT denotes the use of the Full Table for end-to-end evaluation, while MT represents the Mini-Table variant.
Unlike end-to-end trained baselines that shift most computation to offline training and then rely on a single, fully embedded corpus at inference time, CRAFT’s cascaded architecture strategically concentrates expensive computation only where it provides the greatest benefit. This modular design offers two key advantages: (1) it avoids costly retraining when adapting to new domains, and (2) it enables seamless integration of emerging embedding models by simply swapping components without architectural changes.
Retrieval Efficiency.
CRAFT shifts the bulk of computation to an offline pre-computation phase, drastically reducing online embedding calls. We build a SPLADE inverted index over all tables using sparse, lightweight lexical representations and pre-encode each table row with the Stage 2 dense ranker model exactly once. At inference, CRAFT executes only three query embeddings plus 5,100 mini-table embeddings per query:
  • Stage 1 (Sparse Retrieval): One query encoding + inverted-index lookup over 169,898 (NQ-Tables) / 419,183 (OTT-QA) tables.
  • Stage 2 (Dense Pruning): One all-mpnet-base-v2 query embedding + sublinear-time ANN search over all precomputed row embeddings to select the top 5,000 mini-tables, avoiding the $\mathcal{O}(N)$ cost of brute-force search.
  • Stage 3 (Re-ranking): One text-embedding-3-small query embedding + embeddings for the top 100 mini-tables.